### v2

Abstract
Ghost imaging (GI) has demonstrated diverse imaging capabilities arising from its encoding-decoding-based computational imaging mechanism. Accordingly, information-theoretic studies have emerged as a promising direction for characterizing the fundamental performance limits of GI and related computational imaging paradigms. Within the information-theoretic framework, the fundamental question of information-optimal encoding design in GI remains largely unexplored, primarily due to the intractability of the prior probability density function (PDF) of an unknown scene. Here, by leveraging the ability of recursively estimating the PDF of the object to be imaged via Bayesian filtering, we establish an adaptive information-maximization encoding (AIME) design framework. Based on the adaptively estimated posterior PDF from previously acquired measurements, the expected information gain of subsequent detections is evaluated and maximized to design the corresponding encoding patterns in a closed-loop manner. Under this framework, the theoretical form of the information-optimal encoding under the energy-constrained regime is analytically derived. Corresponding experimental results show that, GI systems employing information-optimal encoding achieve markedly improved imaging performance compared with conventional fixed point-to-point imaging without relying on additional heuristic regularization, particularly in low signal-to-noise ratio regimes. Moreover, the proposed AIME strategy is compared with existing encoding strategies under a representative experimental constraint, which consistently shows that it enables significantly enhanced information acquisition capability compared with existing encoding strategies, leading to substantially improved imaging quality. These results establish a principled information-theoretic foundation for optimal encoding design in computational imaging paradigms, provided that the forward model can be accurately characterized.
1 Introduction
Ghost imaging (GI) [moreau2018ghost] has demonstrated a broad range of capabilities, including acquisition of high-dimensional (H-D) information through low-dimensional detection [eshun20253d], enhancement of two-dimensional resolution via H-D light-field information [tong2020breaking, tong2025single], non-locally decoupling detection efficiency and imaging resolution [yuan2021single], and reduction of the unnecessary detection redundancy [hu2019optimization]. These advantages have enabled its diverse applications in imaging LiDAR [deng2017performance, pan2021micro], single-shot H-D imaging cameras [liu2020spectral, chu2021spectral], super-resolution microscopy [li2019single, chen2025multicolor], low-dose and multi-dimensional radiographic imaging [zhang2018tabletop, he2021single], imaging in special band [stantchev2020real, bogdanov2025ghost], and task-oriented imaging [wu2025image, abbas2025target]. From the perspective of imaging information acquisition [barrett2013foundations, shensheng2022review], GI can be interpreted as an encoding-decoding-based computational imaging paradigm [hu2022ghost], where the imaging object is encoded by randomly fluctuating light fields, and the desired image is decoded through high-order light-field correlation. This formulation naturally places GI within an information-theoretic framework, under which analyses of information acquisition efficiency and fundamental performance are in principle attainable [li2017negative, hu2021correspondence, zhou2024theoretical].
Within the information-theoretic framework, a fundamental question in GI, and more broadly in such computational imaging paradigms, concerns the existence and design of information-theoretically optimal encoding [wei2016mutual]. In information theory, the encoding optimality is commonly evaluated using information-content measures, such as mutual information [blahut1987principles] and Fisher information [fisher1922mathematical]. However, the probability density function (PDF) of the object to be imaged, which is required for calculating these information measures, is generally intractable in practical imaging systems. As a result, information-theoretic studies on encoding design of GI have so far been confined to very restrictive assumptions or special scenarios [enrong2013mutual] rather than a general imaging scenario. As a practical compromise, encoding optimization has been extensively explored within the framework of compressed sensing (CS), motivated by the principle that incoherent sampling can preserve the information of sparse signals with high efficiency [eldar2012compressed]. Assuming sparsity of the image in a specific transform domain, the encoding patterns in GI can be optimized by minimizing mutual coherence of the sensing matrix [xu2015optimization, czajkowski2018single, hu2019optimization]. Nevertheless, such CS-based framework typically neglect the influence of measurement noise, as exact recovery guarantees are theoretically established only in the noiseless case [tropp2004greed]. In practical scenarios, the detection noise is often amplified [arias2011noise] in related image retrieval, rendering signal recovery particularly challenging in the low signal-to-noise ratio (SNR) regime. Thus, existing approaches along this line remain theoretically incomplete, and are difficult to apply to GI under realistic low-SNR conditions.
Recently, by modeling the linear computational imaging system within the Bayesian filtering framework, Du et al. [du2025information] proposed a recursive approach to estimate the PDF of the imaging object from already-acquired detection signals. This formulation suggests a promising route for evaluating information-theoretic measures and designing information-optimal encoding light fields in an adaptive manner. However, from existing studies on adaptive GI [wu2020online, wang2023dual], it remains unclear whether adaptive measurements can consistently yield fundamental performance gains, particularly from an information-theoretic perspective. More broadly, in the context of adaptive CS, the superiority of adaptive sensing over non-adaptive measurements has long remained theoretically inconclusive [donoho2006compressed, arias2012fundamental, malloy2014near], especially in noisy and practically constrained regimes.
In this Article, we explicitly investigate the possibility to adaptively establish the information-optimal encoding for GI, and propose a general adaptive information-maximization encoding (AIME) framework. It enables quantitative evaluation of the information content acquired in the encoding-detection process, and facilitates the design of optimal encoding under practical constraints through information maximization. Theoretical analysis and experimental results consistently reveal that the form of information-optimal encoding is intrinsically governed by the underlying physical constraints. Under an ideal total-energy constraint corresponding to fixed illumination energy, the information-maximization encoding converges to a point-like illumination pattern, leading to an adaptive scanning strategy whose order is determined by the evolving object statistics rather than by a predetermined sequence. Owing to this adaptive design, AIME achieves superior imaging performance compared to conventional “fixed point-to-point” imaging modes without relying on additional heuristic regularization, particularly in low SNR regimes. Under realistic modulation constraints imposed by experimental hardware, although the optimal encoding no longer reduces to a point-like structure, it still emerges as an object-dependent adaptive pattern. In this regime, comparisons with representative encoding strategies show that AIME consistently provides improved imaging quality across a wide range of sampling ratio (SR) and SNR conditions. Further evaluation of the acquired information of different encoding strategies shows strong agreement with image quality, indicating that the performance gain originates from more efficient information acquisition rather than heuristic estimation effects. The proposed framework establishes a principled information-theoretic foundation for characterizing and approaching the fundamental performance of computational imaging paradigms.
2 Results
2.1 Principle of AIME for GI
We formulate AIME as a closed-loop information-theoretic framework for GI, in which the encoding patterns are treated as variables to be optimized according to the evolving statistical state of the estimated imaging object. As illustrated in Fig. 1 b, at each measurement step, the statistical state is represented by a posterior PDF inferred from previously detected signals, based on which the expected information gain of a candidate encoding is evaluated. The subsequent encoding is then determined by maximizing the information criterion under explicit physical followed by a recursive update of the posterior PDF after measurement. Within this formulation, encoding, detection, and statistical inference in GI are integrated into a unified framework, enabling information-optimal encoding to be defined and implemented sequentially without requiring special assumption on prior knowledge of the object statistics.
2.1.1 Bayesian formulation for recursive posterior PDF tracking
To enable information-theoretic evaluation and optimization of encoding, the statistical state of the imaging object is characterized by a posterior PDF that is updated recursively as measurements are acquired. To this end, the forward encoding-detection process of GI is formulated to construct the measurement-likelihood function $p({\bf z}\mid{\bf x})$, specifically as [hu2022ghost, liu2025comprehensive]
${\bf z}={\beta\bf H}{\bf x}+{\bf n},$ (1)
where $\bf z$} stands for the recorded signal, $\bf x$} denotes the desired object’s image, $\bf H$} is the measurement matrix consisting of the encoding light fields $\left\{{\bf h}_{1},{\bf h}_{2},\cdots,{\bf h}_{K}\right\}^{\top}$ with $K$ being the number of acquired measurements, $\bf n$ is the detection noise, and $\beta$ represents a coefficient accounting for various system-dependent factors. This formulation is also shared by a broad class of computational imaging systems.
Then, an initial Gaussian distribution $p({\bf x}\mid{\bf Z}_{0})=p_{0}({\bf x})={\mathcal{N}}(\widehat{\bf x}_{0},\widehat{\bf P}_{0})$ is assumed by incorporating statistics of imaging scenes in the frequency domain [torralba2003statistics](see Methods and Materials for more details). By considering the forward sub-model with a frame of light fields ${\bf h}_{k}$ constructed as $p({\bf z}_{k}\mid{\bf x})={\mathcal{N}}(\beta{\bf h}_{k}^{\top}{\bf x},R_{k})$, the posterior PDF after the $k$-th measurement is obtained as
$p({\bf x}\mid{\bf Z}_{k})=\mathcal{N}(\widehat{\bf x}_{k},\widehat{\bf P}_{k}).$ (2)
Here a Gaussian approximation $n_{k}\sim{\mathcal{N}}(0,R_{k})$ for the noise model is adopted, where $R_{k}=\omega^{2}{\bf h}_{k}^{\top}\widehat{\bf x}_{k-1}$ results from the photon noise assumption [liu2025comprehensive], with the current estimate $\widehat{\bf x}_{k-1}$ used to approximate the unknown true $\bf x$}. Under these settings, the posterior statistics ${\widehat{\bf x}}_{k}$, $\widehat{\bf P}_{k}$ are obtained via recursive expressions
${{\widehat{\bf x}}_{k}}={\widehat{\bf x}}_{k-1}+{{\bf K}_{k}}({{\bf z}_{k}}-\beta{{\bf h}_{k}^{\top}}{\widehat{\bf x}_{k-1}}),$ (3a)
$\widehat{\bf P}_{k}=({\bf I}-\beta{{\bf K}_{k}}{{\bf h}_{k}^{\top}})\widehat{\bf P}_{k-1},$ (3b)
with
${{\bf K}_{k}}=\beta\widehat{\bf P}_{k-1}{\bf h}_{k}{(\beta^{2}{{\bf h}_{k}^{\top}}\widehat{\bf P}_{k-1}{\bf h}_{k}+{R_{k}})^{-1}}.$
These posterior statistics serve as sufficient descriptors of the statistical state and provide the basis for quantitatively assessing the information content of subsequent measurements, as will be used in the following subsection.
Figure 1: a. Schematic illustration of the ghost imaging (GI) system based on a spatial light modulator. b. Overview of the proposed adaptive information-maximization encoding (AIME) framework for GI. The optimal encoding is designed by maximizing the information contained in the detection signal of the object to be imaged. The information is calculated based on the forward detection model and posterior PDFs.
2.1.2 Theoretical objective of AIME
Based on the posterior PDF $p({\bf x}\mid{\bf Z}_{k})$ obtained from recursive inference, the encoding for the $(k+1)$-th measurement can be formulated as a variable to be optimized according to an information-theoretic criterion. In the AIME framework, this optimization is guided by the expected reduction of posterior uncertainty induced by a candidate encoding, which provides a principled definition of information-optimal encoding. Here, the mutual information and the Fisher-information-related criterion are employed as two representative information-theoretic measures of this principle.
  • The mutual information between the $k$-th detection signal and the imaging object ${\rm I}(z_{k},{\bf x}\mid{\bf Z}_{k-1})$ is calculated by
    ${\rm I}({\bf z}_{k},{\bf x}\mid{\bf Z}_{k-1})=\int{\rm d}{\bf z}_{k}{\rm d}{\bf x}p({\bf z}_{k},{\bf x}\mid{\bf Z}_{k-1}))\log\frac{p({\bf z}_{k},{\bf x}\mid{\bf Z}_{k-1})}{p({\bf z}_{k})p({\bf x}\mid{\bf Z}_{k-1})}$ (4)
    and obtained as
    ${\rm I}({\bf z}_{k},{\bf x}\mid{\bf Z}_{k-1})=\frac{1}{2}\log\frac{1}{\left|{\bf I}-\beta{{\bf K}_{k}}{{\bf h}_{k}^{\top}}\right|}=\frac{1}{2}\log\left(1+\beta^{2}\frac{{{{\bf{h}}_{k}^{\top}}{{\widehat{\bf{P}}}_{k-1}}{\bf{h}}_{k}}}{{{R_{k}}}}\right).$ (5)
    Hence, maximizing the mutual information is equivalent to maximizing the following objective
    ${\mathcal{L}}_{\rm MI}=\frac{{{{\bf{h}}_{k}^{\top}}{{\widehat{\bf{P}}}_{k-1}}{\bf{h}}_{k}}}{{{R_{k}}}}.$ (6)
    For convenience, this mutual-information-based formulation is hereafter referred to as AIME-MI.
  • The Fisher information contained in the detection signal $z_{k}$ is in fact represented by the Fisher information matrix (FIM) ${\widehat{\bf P}}_{k}^{-1}$ [du2025information]. Since the calculation of matrix inverse is time-consuming, instead of FIM, here the Cramér-Rao bound (CRB) is applied to perform the information-optimal encoding design by minimizing it. Specifically, the mean CRB (mCRB) which stand for the error bound averaging on all pixels of the imaging object is calculated by
    ${\rm mCRB}={\rm Tr}({\widehat{\bf P}}_{k})$ (7)
    and obtain as
    ${\rm mCRB}={\rm Tr}({\widehat{\bf P}}_{k-1})-\frac{{\bf h}_{k}^{\top}\widehat{{\bf P}}_{k-1}^{2}{\bf h}_{k}}{{\bf h}_{k}^{\top}\widehat{{\bf P}}_{k-1}{\bf h}_{k}+\frac{R_{k}}{\beta^{2}}}.$ (8)
    Since the covariance matrix $\widehat{\bf{P}}_{k-1}$ has been determined by already-acquired detection, to minimize the mean CRB, only the last term in Eq. (8) can be optimized. Hence, minimizing the mCRB is equivalent to maximizing the following objective
    ${\mathcal{L}}_{\rm CRB}=\frac{{\bf h}_{k}^{\top}\widehat{{\bf P}}_{k-1}^{2}{\bf h}_{k}}{{\bf h}_{k}^{\top}\widehat{{\bf P}}_{k-1}{\bf h}_{k}+\frac{R_{k}}{\beta^{2}}}.$ (9)
    For convenience, this formulation is hereafter referred to as AIME-CRB.
With Eqs. (6) and (9), it is possible to optimize the subsequent encoding ${\bf h}_{k}$ by
$\max_{{\bf h}_{k}}{\mathcal{L}},~~~{\rm s.t.}~~{\mathcal{C}}({\bf h}_{k})$ (10)
where ${\mathcal{C}}({\bf h}_{k})$ denotes the practical constraints on the encoding ${\bf H}_{k}$. We would like to note that, although derived from an estimation-theoretic perspective, the CRB-based criterion quantifies the same objective of posterior uncertainty reduction as the mutual-information-based formulation.
Under this formulation, the design of information-optimal encoding is unified as a constrained optimization problem driven by posterior PDF. Once a specific information criterion and the physical constraints on the encoding are specified, the structure of the optimal encoding follows directly from the solution of this optimization problem. Consequently, the AIME framework itself is not tied to a particular form of information measure or constraint, but provides a general principle for encoding design that can be instantiated under different system configurations.
2.2 Information-optimal encoding in an ideal energy-constrained regime
Figure 2: Illustration of the theoretical result of AIME-$\ell_{1}$ patterns, each pattern is normalized to its maximum value for visualization. Figures a1, a2 and b1, b2 respectively show AIME results of two different imaging objects. a1 and b1 are patterns obtained by analytical solution of Eq. (12), each of which show exactly a scanning point; and the position of the point is marked with a purple arrow and box. a2 and b2 are patterns obtained by minimizing the mean CRB, which gradually become a small-size spot as the number of measurements increases; and the spot region in each pattern is displayed in an enlarged view. The first Hermite Gaussian pattern is employed as the first frame of encoding pattern here for all cases.
Based on the information-theoretic objective derived in the previous subsection, we first examine AIME in an idealized yet physically meaningful encoding regime. Specifically, we consider a total-energy constraint on the encoding light fields, which captures a fundamental limitation in optical imaging systems. Since the total energy of encoding ${\bf h}_{k}$ can be represented by its $\ell_{1}$ norm, i.e., $\sum_{j}{{\bf h}_{k}}_{j}=\|{\bf h}_{k}\|_{1}$, Aime under this constraint is denoted with suffix $\ell_{1}$”. In this regime, the optimal encoding admits a closed-form analytical solution, as established in the following theorem.
Theorem 1 (Theoretical solution of information-optimal encoding under a total-energy constraint).
Consider the following optimization problem:
$\max_{\mathbf{h}_{k}}$ $\frac{\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}}{\mathbf{h}_{k}^{\top}\mathbf{x}},$ (11)
$\rm{s.t.}$ $\left\|\mathbf{h}_{k}\right\|_{1}=C,\mathbf{h}_{k}\succeq\mathbf{0},$
where $\mathbf{P}_{k-1}$ is a positive definite covariance matrix, $\mathbf{x}$ is a non-negative vector, and $C>0$ is a constant. Then the above problem admits an analytical optimal solution. Specifically, the optimal encoding vector is given by
$\hat{\mathbf{h}}_{k}=\mathbf{e}_{\hat{i}},\quad\hat{i}=\arg\max_{i}\frac{\left(\mathbf{P}_{k-1}\right)_{ii}}{x_{i}},$ (12)
where $\mathbf{e}_{\hat{i}}$ denotes the standard basis vector whose $\hat{i}-$th element is one and all other elements are zero.
This Theorem establishes that, in the ideal energy-constrained regime, information-optimal encoding necessarily reduces to an adaptive point-wise scanning strategy. Importantly, this point-wise structure does not imply a predetermined or fixed scanning order, but instead, varies according to the evolving posterior statistics. For the CRB-based objective shown as Eq. (9), although a closed-form analytical solution of the optimal encoding is difficult to derive, simulation and experimental results indicate a consistent structural evolution behavior: the optimal encoding initially appears as a diffuse illumination pattern and gradually concentrates into a localized spot as the number of acquired measurements increases.
Figure 3: Experimental comparison of imaging results of GI with AIME-$\ell_{1}$ and conventional imaging. a: Representative imaging results obtained by GI AIME-$\ell_{1}$ and conventional imaging under sampling ratios (SRs) of 50% and 100%. Results are shown for two signal-to-noise ratio (SNR) levels, as indicated on the left. The reference image is shown on the right, which is obtained by full-sampling point-wise scanning under a significantly high SNR. b: Quantitative evaluation of imaging quality in terms of PSNR and SSIM as functions of the sampling ratio for the case of ${\rm SNR}=7.03~{\rm dB}$. The horizontal gray lines indicate the performance of conventional full-sampling (100% SR) imaging for comparison. c: Corresponding PSNR and SSIM curves for the case of ${\rm SNR}=4.85~{\rm dB}$. The horizontal gray lines again indicate the performance of conventional full-sampling (100% SR) imaging for comparison.
Figure 2 provides an illustrative visualization of the AIME-$\ell_{1}$ encoding patterns, where panels a and b correspond to two representative imaging objects. To ensure a fair comparison, the first encoding pattern is identically initialized as the first Hermite Gaussian mode [huang2024compressed] for all cases. The results visually corroborate the theoretical prediction that information-optimal encoding evolves toward an adaptive point-wise structure under the total-energy constraint, which fundamentally differs from conventional raster-scan imaging in that the scanning order is not fixed but adaptively determined by the statistical structure of the object. This difference between AIME-$\ell_{1}$ and conventional imaging lead to an improved imaging performance. To show that, A quantitative experimental comparison under identical total energy and noise conditions is performed. As shown in Fig. 3, the experimental results indicate that AIME consistently achieves higher imaging quality than traditional imaging, particularly in low-SNR regimes. In specific, with approximately 50% sampling, AIME yields a visually recognizable imaging result at a detection SNR of 7.03 dB, whereas conventional imaging with 100% sampling fails to recover the object reliably. When both methods operate at full sampling, AIME produces noticeably improved image quality, with an increase of about 0.2 in SSIM. As the detection SNR further decreases to 4.85 dB, the object becomes nearly indistinguishable in conventional imaging, while AIME remains capable of providing a coarse yet informative image. Details of the experimental setting for this comparison is presented in Materials and Methods, and more experimental comparison results on representative scenes are provided in the supplementary materials. These results demonstrate that the AIME strategy enables more efficient information utilization than fixed raster-scan acquisition in the ideal energy-constrained regime.
An alternative intuition for the emergence of point-wise optimal encoding can be drawn from CS. It is well established in CS theory that sparse signals can be efficiently acquired through incoherent measurements in a conjugate domain, for example, spatially sparse signals can be recovered from partial Fourier measurements [candes2006robust]. In the present imaging scenario, natural scenes are often sparse in transform domains such as the discrete cosine transform (DCT) domain. From this perspective, concentrating measurements on a limited set of spatial locations can be viewed as an intuitive counterpart to selectively probing sparse degrees of freedom in a conjugate domain, offering a qualitative analogy to the point-wise scanning behavior predicted by the information-theoretic analysis.
2.3 Information-optimal encoding under realistic modulation constraints
The preceding analysis establishes the intrinsic structure of information-optimal encoding under an idealized total-energy constraint. In practical GI systems, however, additional modulation constraints on the encoding patterns are inevitably imposed. It is therefore essential to examine whether the advantages of AIME persist under such realistic conditions. To this end, beyond the ideal total-energy constraint, we further impose a bounded-amplitude constraint on the encoding patterns,
$0\leq h_{k,i}\leq C,\quad\forall i,$ (13)
which reflects realistic hardware limitations such as the finite modulation depth of spatial light modulators or digital micromirror devices. To differentiate with $\ell_{1}, AIME$ under this constraint is denoted with suffix $\ell_{\infty}$”, since the maximal element of non-negative encoding ${\bf h}_{k}$ can be represented by its $\ell_{\infty}$ norm.
Under such constraints, the closed-form analytical solution derived above no longer applies, and the information-optimal encoding is therefore obtained numerically by solving the resulting constrained optimization problem. Figure 4 provides an illustrative visualization of the joint evolution of imaging results and the optimized encoding patterns under the proposed strategy. Two different imaging scenes, namely cameraman and pepper, are considered, and the peak signal-to-noise ratio (PSNR) is used to quantify the imaging quality. Figures 4 (a1) and (b1) show imaging results together with the corresponding optimized encoding patterns at different sampling ratios (SRs), where the selected SRs correspond to the marked points in Figs. 4 (a2) and (b2), respectively. The optimized encoding patterns exhibit a clear coarse-to-fine evolution as the SR increases. At low SRs, the patterns consist of large-scale speckle-like structures, whereas at higher SRs, finer spatial features gradually emerge. This behavior is consistent with the gradual refinement of the posterior distribution in the Bayesian filtering process, where early measurements focus on coarse information content and subsequent measurements concentrate on increasingly localized spatial details. The corresponding PSNR curves shown in Figs. 4 (a2) and (b2) demonstrate a rapid improvement in reconstruction quality at the early stage of sampling, indicating the high information efficiency of the proposed strategy under limited measurements. To further illustrate the accuracy and uncertainty evolution of the reconstruction, the posterior probability density functions (PDFs) of two representative pixel locations (marked by red and blue crosses in the estimated images) are plotted in Figs. 4 (a3) and (b3). As SR increases, the posterior PDFs become progressively sharper, reflecting the systematic reduction of estimation uncertainty and the convergence of pixel-wise intensity estimates. By comparing the results obtained for different imaging scenes, it is evident that the optimized encoding patterns are also strongly scene-dependent under the bounded-amplitude constraint. This observation is fully consistent with the theoretical formulation of AIME and highlights its ability to adaptively allocate measurements according to the estimated statistical state of the imaging object.
Figure 4: Simulation results for illustrating the evolution process of the proposed strategy on two different imaging scenes. a1 and b1 show the evolution of estimated images and corresponding optimized encoding patterns with sampling ratio increasing for two imaging scenes, respectively; a2 and b2 show evolution of PSNR indicators with the increase in sampling; a3 and b3 show evolution of estimated PDFs of two selected pixels as marked on the imaging results, respectively.
Figure 5: Experimental imaging results for comparison between AIME-$\ell_{\infty}$ and some typical encoding patterns. Results of AIME-$\ell_{1}$ that obtained under the total-energy constraint are also shown here to serve as a fundamental-limit reference. a1, a2, and a3 show three different representative cases. The SR and SNR for them are: a1 SNR around 18 dB, and 25 % SR; a2 SNR around 18.5 dB, and 37.5 % SR; a3 SNR around 20 dB, and 37.5 % SR. PSNR and SSIM indicators of imaging results corresponding to three scenes as SR increases are shown in b1, c1; b2, c2, and b3, c3, respectively. To further quantify and confirm the information acquisition enhancement, the acquired information of different encoding strategies is calculated, and shown in d1, d2, and d3.
To evaluate the practical effectiveness of AIME-$\ell_{\infty}$ under such realistic experimental constraints, we compare it with several representative existing encoding patterns in GI, including pseudo-thermal random patterns, Walsh-Hadamard patterns, and Hermite Gaussian patterns [huang2024compressed]. For theoretical reference, we also include the ideal total-energy-constrained AIME-$\ell_{1}$ results, which serves as a theoretical performance limit without experimental modulation constraints. All results are obtained by employing the same Bayesian filtering framework to ensure a fair comparison that isolates the effect of encoding design. Figure 5 summarizes the results of three representative scenes acquired under different sampling ratios (SRs) and signal-to-noise ratios (SNRs). The imaging results are shown as Figs. 5 a1-a3, respectively. For all tested scenes, both AIME-based strategies (i.e., AIME-MI and AIMe-CRB) consistently yield superior imaging quality compared with other encoding patterns under the same modulation constraints, as evidenced by clearer structural details and significantly reduced noise artifacts. The corresponding PSNR and SSIM values are reported under each image, which confirm that AIME achieves much better quantitative performance across all conditions. At the same time, a performance gap between AIME-$\ell_{\infty}$ and AIME-$\ell_{1}$ theoretical reference can be observed, suggesting the loss of achievable information imposed by hardware constraints on the encoding space. The evolution of PSNR and SSIM as functions of the SR is further plotted in Figs. 5 b1-b3 and c1-c3, respectively. These results reveal that AIME-$\ell_{\infty}$ exhibits a pronounced advantage at early measurement stages, where substantial imaging quality improvements are already achieved at low SRs. As the number of measurements increases, AIME-$\ell_{\infty}$ continues to provide steady performance gains, whereas the imaging quality obtained with other encoding patterns rapidly saturates. This behavior demonstrates the strong compressed-sampling capability of AIME and its effectiveness in allocating the most informative measurements. The gap between AIME-$\ell_{\infty}$ and AIME-$\ell_{1}$ can be also reflected from imaging quality curves, which quantifies the performance loss induced by realistic modulation constraints.
To further quantify the information acquisition capability of different encoding strategies, we explicitly compute the cumulative mutual information accumulated from obtained detection signals. Specifically, the cumulative mutual information is calculated as $\sum_{k}\frac{1}{2}\ln\left(1+\frac{{{\bf H}_{k}\widehat{\bf P}_{k-1}{\bf H}_{k}^{\top}}}{{R_{k}}}\right)$ and plotted on the logarithmic scale. The corresponding results shown in Figs. 5 d1-d3 represent evolution of the acquired amount of information about the unknown imaging scene during the detection process. It can be found that, across different SRs and noise levels, AIME consistently enables the detection signals to acquire substantially more information about the imaging object than other comparison encoding strategies. Also, the separation between AIME-$\ell_{\infty}$ and AIME-$\ell_{1}$ curves quantifies the information loss due to realistic modulation constraints, and mirrors the gap observed in reconstruction quality. Moreover, the ordering of the information curves agrees with the ordering of reconstruction quality, establishing a direct correspondence between information acquisition and imaging performance. This agreement confirms that the superiority of AIME arises from principled information maximization rather than other heuristic effects.
3 Discussion
In this work, an adaptive information-maximization encoding strategy is developed for GI. The proposed framework is intended to serve as a generally applicable approach for information-optimal encoding in computational imaging systems with the encoding-decoding mode, provided that the forward imaging model can be explicitly characterized. In the present study, the forward model is formulated as a simplified linear model and demonstrated using a spatial-domain GI system with sequential bucket detection. While this simplified model enables analytical tractability and conceptual clarity, more realistic forward models accounting for optical blurring, down-sampling, and noise [huang2020ghost, huang2021ghost, liu2025comprehensive] can be further incorporated into the proposed framework. Moreover, since Bayesian filtering theory is not restricted to linear systems, the proposed strategy can, in principle, be generalized to computational imaging problems governed by nonlinear forward models, such as the coherent detection GI [pan2021micro]. Beyond the form of the forward model, the framework is also not restricted to a specific sensing configuration. Although AIME is demonstrated with sequential bucket detection, the underlying Bayesian filtering and information optimization principles can be naturally extended to batch-measurement settings. In such cases, multiple measurements are acquired simultaneously, and the framework can be generalized to adaptive or semi-adaptive encoding design in a higher-dimensional measurement space. This suggests its applicability to a broader class of computational imaging systems with array detection, like the GI camera [chu2021spectral].
In addition to the forward model, another essential component of the proposed framework is the setting of the initial image PDF. In the presented study, a general prior is adopted by imposing constraints in the image frequency domain, which allows an PDF to be defined for an entirely unknown imaging scene. While this choice provides a concise and robust prior for Bayesian filtering, it can be further improved by incorporating alternative analytical image prior that better capture higher-order statistical properties of natural images [hyvarinen2009natural]. Furthermore, with the development of deep neural models, image priors can now be effectively learned from data using deep generative models. In scenarios where a sufficiently large and representative training dataset is available, the proposed strategy can be naturally combined with image priors provided by trained generative models [elata2024adaptive, pinkard2024information]. Such learned priors are expected to offer a more expressive statistical representation than the general analytical prior employed in this work, potentially leading to further performance improvements.
Altogether, this work highlights the potential of information-theoretic encoding design as a principled approach to GI and, more broadly, to computational imaging systems. By explicitly modeling the information carried by detection signals under realistic physical constraints imposed by the imaging system, the proposed framework provides a systematic methodology to characterize and approach fundamental imaging limits. The emergence of point-like encoding as information-optimal under a total-energy constraint offers an information-theoretic insight suggesting that point-to-point sampling strategies observed in biological vision [land2012animal] may reflect a principle of efficient information acquisition. Beyond GI, the proposed framework is expected to inspire further information-theoretic studies on the evaluation and optimization of imaging and sensing systems, with potential implications for a broad range of applications, including autonomous sensing, machine vision, and data-efficient perception tasks. Importantly, in numerous such systems, flexible light-field modulation platforms such as metasurfaces and programmable photonic devices are increasingly employed as core encoding components, which provides structured design space for encoding optimization. Incorporating this into the proposed framework is supposed to offer a principled pathway that integrates emerging hardware platforms with the information-optimal encoding design.
Materials and Methods
Details for encoding optimization
The encoding optimization problem is formulated as a constrained optimization task and solved using Projected Gradient Descent (PGD), which alternates between gra- dient descent updates and projection onto the feasible constraint set. During the gradient descent step, we employ the SGD optimizer with learning rates of $1\times 10^{9}$ and $1\times 10^{5}$ for the mutual-information-based and CRB-based objectives, respectively. All remaining optimizer hyperparameters follow the default PyTorch configuration. To balance computational efficiency and estimation accuracy, each encoding optimization is performed using 100 PGD iterations. All numerical calculation are implemented in PyTorch 2.5.1 with CUDA 12.1 and executed on an NVIDIA GeForce RTX 3090 GPU. For a $128\times 128$ image, one complete cycle consisting of Bayesian filtering and encoding optimization requires approximately 0.12 s.
Measurement of parameters $\beta$ and $R_{k}$ in the experiment
The experiment is performed with a GI system with DMD modulation. To determine the forward model in Eq. (1), the scaling factor $\beta$ and the noise statistics $\{R_{k}\}$ are experimentally calibrated prior to imaging.
Calibration of $\beta$
A whiteboard with uniform reflectance is used as the calibration object. A set of $M$ Hadamard encoding patterns ${\bf h}^{(1)},\cdots,{\bf h}^{(M)}$ is projected sequentially, and for each pattern $L$ repeated measurements are recorded, yielding detection vectors ${\bf z}^{(m)}\in{\mathcal{R}}^{L}$. Then, $\beta$ is estimated by
$\beta=\left\langle\frac{\overline{z}^{(m)}}{\sum_{i}{\bf h}^{(m)}_{i}}\right\rangle_{m}$ (14)
where $\overline{z}^{(m)}$ denotes the average on elements of ${\bf z}^{(m)}$ and $\left\langle\cdot\right\rangle_{m}$ represents averaging over all calibration encoding patterns.
Calibration of noise statistics
As modeled in Section 2.1.1, the noise variance is parameterized as
$R_{k}=\omega^{2}{\bf h}_{k}^{\top}{\bf x}$
where $\omega^{2}$ characterizes the proportionality between signal intensity and noise variance and therefore controls the effective detection SNR. With those measurements on the whiteboard by using Hadamard patterns, estimation on $\omega^{2}$ in practical conditions is as follows. For the $m$-th encoding pattern, with $L$ repeated measurements, the corresponding noise coefficient is estimated as
${\omega^{2}}^{(m)}=\frac{{\sigma^{2}}^{(m)}}{{\bar{z}}^{(m)}}$ (15)
where ${\sigma^{2}}^{(m)}=\frac{\sum_{l=1}^{L}\left(z_{l}^{(m)}-\bar{z}^{(m)}\right)^{2}}{L}$ denotes the sample variance of the detection signal. The calibrated noise parameter is then taken as the ensemble average
$\omega^{2}_{\text{eff}}=\left\langle{\omega^{2}}^{(m)}\right\rangle_{m}.$ (16)
By adjusting the illumination intensity and repeating the calibration procedure, noise parameters corresponding to different operating SNR regimes are obtained. In each regime, the calibrated $\omega^{2}_{\text{eff}}$ is used consistently in both mutual-information-based and CRB-based encoding optimization.
Experimental setting in comparison between GI with AIME-$\ell_{1}$ and conventional imaging
For experimental comparison between GI with AIME-$\ell_{1}$ and “fixed point-to-point” conventional imaging, both strategies are implemented using the same DMD-based GI system to ensure identical optical and detection conditions. For AIME-$\ell_{1}$, the optimized encoding patterns are physically realized with DMD. To improve illumination stability, each $4\times 4$ block of DMD micro-mirrors is grouped to form a single effective illumination pixel. The optimized patterns are mapped onto these grouped pixels and sequentially projected onto the scene, and the corresponding bucket detection signals are recorded. For conventional imaging, the same DMD-based GI system is operated in a raster-scanning mode. Specifically, a sequence of point-like illumination patterns is loaded onto the DMD, where each grouped $4\times 4$ block is illuminated one at a time following a fixed scanning order. The detected bucket signal associated with each illumination position is directly assigned as the grayscale value of the corresponding image pixel, thereby forming the image.
In the experiment under low-SNR condition, additional background noise is introduced using scattering light generated by a rotating ground glass. This produces a dynamic background whose temporal fluctuation, after bucket detection and time averaging, can be well approximated as additive Gaussian noise with constant variance. To quantify the noise level, $L$ measurements are acquired using a completely dark pattern. Let $z_{\text{back}}^{(l)}$ denote the $l$-th recorded signal, and $\bar{z}_{\text{back}}$ its sample mean over $L$ measurements, the background noise variance $\sigma_{\text{back}}^{2}$ is estimated as:
$\sigma_{\text{back}}^{2}=\frac{\sum_{l=1}^{L}\left(z_{\text{back}}^{(l)}-\bar{z}_{\text{back}}\right)^{2}}{L-1},$ (17)
With experimental detection signal $\{z_{k}\}$ corresponding to both AIME and raster-scanning, the experimental SNR is estimated by
$\text{SNR}=10\log_{10}\left(\frac{{\bar{z}}-\bar{z}_{\text{back}}}{\sigma_{\text{back}}}\right).$ (18)
where $\bar{z}$ is the average intensity of detection signals.
Setting the initial image PDF
In the proposed strategy, an initial Gaussian PDF $p({\bf x})={\mathcal{N}}({\bf x}_{0},{\bf P}_{0})$ of the object’s image is set. To specify the parameters ${\bf x}_{0}$ and ${\bf P}_{0}$,we adopt a commonly used statistical model for natural images, in which the frequency-domain coefficients $\left\{a({\bf f})\right\}$ of images [torralba2003statistics] follow a Laplacian distribution
$p(a({\bf f}))=\frac{1}{2\lambda({\bf f})}\exp\left[-\frac{\left|a({\bf f})-\mu({\bf f})\right|}{\lambda({\bf f})}\right],$ (19)
where $\mu({\bf f})$ for those non-zero-frequency components is set as 0, $\mu({\bf f}_{0})=\mu_{0}$ representing the zero-frequency amplitude is estimated with a uniform illumination, and $\lambda({\bf f})\propto(|{\bf f}|)^{-\gamma}$ with $\gamma\in[1,2]$. It should be noted that, owing to the high dimensonality of image signals and the central limit theorem, although the frequency-domain coefficients are modeled as independent Laplacian variables, the image-domain pixels, being weighted superpositions of a large number of such coefficients, can be well approximated to be subject to a Gaussian distribution. Based on this, the parameters of the initial Gaussian PDF are set as
${\bf x}_{0}=\mu_{0}{\bf 1}$ (20a)
${\bf P}_{0}=2\Psi^{\top}{\bf\Sigma}^{2}\Psi$ (20b)
where $\bf 1$ is an all-$1$ vector, $\bf\Sigma$} is a diagonal matrix whose diagonal elements consist of $\left\{\lambda({\bf f}_{0}),\lambda({\bf f}_{1}),\cdots\right\}$ and $\bf\Psi$ denotes DCT matrix.
Data availability
Data are available from the corresponding author on request.
Funding
This work is supported by the National Natural Science Foundation of China (62201165, 62475270) and the National Key Research and Development Program of China (2024YFF0505601).
Author contributions
C. Hu and S. Han conceived the idea and supervised the study. J. Sun and C. Hu performed most of the theoretical derivation. J. Sun and C. Hu conducted the experiment under discussion with Z. Bo, Z. Liu and M. Chen. L. Du and W. Liu contribute to code modification for simulation and experimental data analyses. J. Sun and C. Hu analyzed the experimental data, and wrote the original draft. All authors discussed the results and contributed to editing and revising the paper.
Competing interests
The authors declare no competing interests.
Appendix A Theoretical derivation of the objective function for AIME-MI and AIME-CRB
In the proposed strategy, the mutual information ${\rm I}(z_{k},{\bf x}\mid{\bf Z}_{k-1})$ is specifically calculated as follows
${\rm{I}}\left({{z_{k}},{\bf{x}}\left|{{{\bf{Z}}_{k-1}}}\right.}\right)=$ $~{\rm H}({\bf x}\mid{\bf Z}_{k-1})-{\rm H}({\bf x}\mid{\bf Z}_{k})$ (21)
$=$ $~\left[\frac{N}{2}(1+\log 2\pi)+\frac{1}{2}\log\left|\widehat{\bf{P}}_{k-1}\right|\right]-\left[\frac{N}{2}(1+\log 2\pi)+\frac{1}{2}\log\left|\widehat{\bf{P}}_{k}\right|\right]$
$=$ $~\frac{1}{2}\log\frac{\left|\widehat{\bf{P}}_{k-1}\right|}{\left|\widehat{\bf{P}}_{k}\right|}=\frac{1}{2}\log\frac{\left|\widehat{\bf{P}}_{k-1}\right|}{\left|\left(\boldsymbol{I}-\beta{\bf{K}}_{k}{\bf{h}}_{k}^{\top}\right)\widehat{\bf{P}}_{k-1}\right|}$
$=$ $~\frac{1}{2}\log\frac{1}{{\left|{{\bf{I}}-\beta{{\bf{K}}_{k}}{{\bf{h}}_{k}^{\top}}}\right|}}.$
where ${\rm H}(\cdot)$ is the information entropy, which can be expressed as $\frac{1}{2}\log{(2\pi e)^{N}}\det\sum$ for a $N$-variable Gaussian distribution, $\sum$ is the covariance matrix of the corresponding variables, and $|{\bf P}|$ denotes the determinant of matrix $\bf P$. By using the property of matrix determinant $\left|{\bf I}+{\bf u}{\bf v}^{\top}\right|=1+{\bf u}^{\top}{\bf v}$, the above expression can be further simplified
${\rm{I}}\left({{z_{k}},{\bf{x}}\mid{{\bf{Z}}_{k-1}}}\right)$ (22)
$=$ $~\frac{1}{2}\log\frac{1}{{1-\beta{\bf{K}}_{k}^{\top}{{\bf{h}}_{k}}}}$
$=$ $~\frac{1}{2}\log\frac{1}{{1-{{\left({{\beta^{2}}{{\bf{h}}_{k}}{{\widehat{\bf{P}}}_{k-1}}{\bf{h}}_{k}^{\top}+{R_{k}}}\right)}^{-1}}{\beta^{2}}{{\bf{h}}_{k}}{{\widehat{\bf{P}}}_{k-1}}{{\bf{h}}_{k}}}}$
$=$ $~\frac{1}{2}\log\frac{1}{{1-\frac{{{\beta^{2}}{{\bf{h}}_{k}}{{\widehat{\bf{P}}}_{k-1}}{\bf{h}}_{k}^{\top}}}{{{\beta^{2}}{{\bf{h}}_{k}}{{\widehat{\bf{P}}}_{k-1}}{\bf{h}}_{k}^{\top}+{R_{k}}}}}}$
$=$ $~\frac{1}{2}\log\frac{{{\beta^{2}}{{\bf{h}}_{k}}{{\widehat{\bf{P}}}_{k-1}}{\bf{h}}_{k}^{\top}+{R_{k}}}}{{{R_{k}}}}$
The mean CRB is calculated as follows
${\rm mCRB}=$ $~{\rm Tr}({\widehat{\bf P}}_{k})$ (23)
$=$ $~\operatorname{Tr}\left(\left(\bf{I}-\beta{\bf{K}}_{k}{\bf{h}}_{k}^{\top}\right)\widehat{\bf{P}}_{k-1}\right)$
$=$ $~\operatorname{Tr}\left(\widehat{\bf{P}}_{k-1}\right)-\operatorname{Tr}\left(\beta{\bf{K}}_{k}{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}\right)$
$=$ $~\operatorname{Tr}\left(\widehat{\bf{P}}_{k-1}\right)-\operatorname{Tr}\left(\beta^{2}\widehat{\bf{P}}_{k-1}{\bf{h}}_{k}\left(\beta^{2}{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}{\bf{h}}_{k}+R_{k}\right)^{-1}{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}\right)$
where ${\rm Tr}({\bf P})$ denotes the trace of matrix $\bf P$. By using the property of matrix trace, the above expression can be rewritten as
${\rm mCRB}=$ $~\operatorname{Tr}\left(\widehat{\bf{P}}_{k-1}\right)-\beta^{2}\operatorname{Tr}\left({\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}^{2}{\bf{h}}_{k}\left(\beta^{2}{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}{\bf{h}}_{k}+R_{k}\right)^{-1}\right)$ (24)
$=$ $~\operatorname{Tr}\left(\widehat{\bf{P}}_{k-1}\right)-\operatorname{Tr}\left(\frac{{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}^{2}{\bf{h}}_{k}}{{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}{\bf{h}}_{k}+\frac{R_{k}}{\beta^{2}}}\right)$
$=$ $~\operatorname{Tr}\left(\widehat{\bf{P}}_{k-1}\right)-\frac{{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}^{2}{\bf{h}}_{k}}{{\bf{h}}_{k}^{\top}\widehat{\bf{P}}_{k-1}{\bf{h}}_{k}+\frac{R_{k}}{\beta^{2}}}$
Appendix B Proof of Theorem 1
Here we give the proof of Theorem 1 in the article, which indicates that the optimal encoding pattern of AIME-MI under the total-energy constraint correspond to a scanning point.
Theorem 2 (Theoretical solution of AIME-MI under the total-energy constraint).
Consider the following optimization problem:
$\max_{\mathbf{h}_{k}}$ $\frac{\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}}{\mathbf{h}_{k}^{\top}\mathbf{x}},$ (25)
$\rm{s.t.}$ $\left\|\mathbf{h}_{k}\right\|_{1}=C,\mathbf{h}_{k}\succeq\mathbf{0},$
where $\mathbf{P}_{k-1}$ is a positive definite covariance matrix, $\mathbf{x}$ is a non-negative vector, and $C>0$ is a constant. Then the above problem admits an analytical optimal solution. Specifically, the optimal encoding vector is given by
$\hat{\mathbf{h}}_{k}=\mathbf{e}_{\hat{i}},\quad\hat{i}=\arg\max_{i}\frac{\left(\mathbf{P}_{k-1}\right)_{ii}}{x_{i}},$ (26)
where $\mathbf{e}_{\hat{i}}$ denotes the standard basis vector whose $\hat{i}-$th element is one and all other elements are zero.
Proof of Theorem 2.
We first note that the objective function is a ratio of two positive homogeneous functions of ${\bf h}_{k}$. Following standard results in fractional programming [schaible1976fractional], maximizing the ratio
$\frac{\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}}{\mathbf{h}_{k}^{\top}\mathbf{x}},$ (27)
is equivalent to finding the largest scalar $\alpha$ such that
$\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}-\alpha\mathbf{h}_{k}^{\top}\mathbf{x}\geq 0$ (28)
admits a feasible solution of ${\bf h}_{k}$ under the same constraints. This lead to an equivalent parametric optimization problem as
$\max_{\mathbf{h}_{k}}\left(\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}-\alpha\mathbf{h}_{k}^{\top}\mathbf{x}\right),\quad\text{ s.t. }\left\|\mathbf{h}_{k}\right\|_{1}=C,~\mathbf{h}_{k}\succeq\mathbf{0},$ (29)
where
$\alpha=\max_{\mathbf{h}_{k}}\frac{\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}}{\mathbf{h}_{k}^{\top}\mathbf{x}}.$ (30)
This can be further rewritten to a standard quadratic form
$\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}-\alpha\mathbf{h}_{k}^{\top}\mathbf{x}=\left(\mathbf{h}_{k}-\frac{\alpha}{2}\mathbf{P}_{k-1}^{-1}\mathbf{x}\right)^{\top}\mathbf{P}_{k-1}\left(\mathbf{h}_{k}-\frac{\alpha}{2}\mathbf{P}_{k-1}^{-1}\mathbf{x}\right)-\frac{\alpha^{2}}{4}\mathbf{x}^{\top}\mathbf{P}_{k-1}^{-1}\mathbf{x}.$ (31)
Since $\mathbf{P}_{k-1}$ is positive definite, the objective is convex with respect to ${\bf h}_{k}$. Under the constraints $\left\|\mathbf{h}_{k}\right\|_{1}=C,$ and $\mathbf{h}_{k}\succeq\mathbf{0}$, the feasible region is a simplex, making the extreme points correspond to scaled standard basis vectors $\left\{C{\bf e}_{i}\right\}$. Therefore, the maximum of the objective function shown as Eq. (31) is attained at one of the extreme points, implying that the optimal solution $\hat{\bf h}_{k}$} correspond to a standard basis vector.
Substituting ${\bf h}_{k}={\bf e}_{i}$ into the original objective yields
$\frac{\mathbf{h}_{k}^{\top}\mathbf{P}_{k-1}\mathbf{h}_{k}}{\mathbf{h}_{k}^{\top}\mathbf{x}}=\frac{\mathbf{e}_{i}^{\top}\mathbf{P}_{k-1}\mathbf{e}_{i}}{\mathbf{e}_{i}^{\top}\mathbf{x}}=\frac{\left(\mathbf{P}_{k-1}\right)_{ii}}{x_{i}}.$ (32)
From this, the optimal index $\hat{i}$ can be obtained by
$\hat{i}=\arg\max_{i}\frac{\left(\mathbf{P}_{k-1}\right)_{ii}}{x_{i}},$ (33)
which leads to the optimal $\hat{\bf h}_{k}={\bf e}_{\hat{i}}$ and complete the proof. ∎
Appendix C Definition of the detection signal-to-noise ratio
Here we would like to discuss about the definition of the detection SNR in GI systems. In the field of GI, the bucket detection SNR (BSNR) [du2025information] is calculated as
${\rm BSNR}=10\log_{10}\left(\frac{\overline{z^{2}}-\overline{z}^{2}}{\sigma^{2}}\right),$ (34)
to measure the level of signal fluctuations. However, this definition only apply to GI systems with randomly fluctuating light fields, and it is not directly related to the detection process.
In the field of signal processing, the SNR is defined as
${\rm SNR}=10\log_{10}\left(\frac{P_{s}}{P_{n}}\right)=10\log_{10}\left(\frac{|z|^{2}}{|n|^{2}}\right),$ (35)
where $P_{s}$ denotes the signal power and $P_{n}$ is the noise power.
Different from traditional signal processing which deal with electronic signals, in GI the signal is in fact an optical one. Since the detector directly measures the intensity of light fields, the recorded intensity signal can already be considered as the power. Hence, in this study we use the definition as [bian2024broadband]
${\rm SNR}=10\log_{10}\left(\frac{\overline{z}}{\sigma}\right),$ (36)
where $\overline{z}$ represents the average value of the detection signal and $\sigma$ denotes the average noise level.
Appendix D Detailed comparison results between GI with AIME-$\ell_{1}$ and fixed point-to-point conventional imaging under different detection SNRs
Figures A6 to A9 show experimental results of four representative imaging scenes.
Figure A6: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $4.85$ dB SNR, b1, b2, b3 are results under $7.03$ dB SNR.
Figure A7: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $3.28$ dB SNR, b1, b2, b3 are results under $5.79$ dB SNR.
Figure A8: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $3.89$ dB SNR, b1, b2, b3 are results under $5.79$ dB SNR.
Figure A9: Comparison between GI with AIME-$\ell_{1}$ and conventional imaging on scene 1 under varying SRs and SNRs, including imaging results and imaging quality indicators. a1, a2, a3 are results under $2.09$ dB SNR, b1, b2, b3 are results under $4.30$ dB SNR.
Appendix E Detailed comparison between GI with AIME and existing encoding patterns
Figures A10 to A13 show experimental results of four representative imaging scenes for comparing GI with AIME and existing encoding patterns.
Figure A10: Experimental imaging results of scene 1 under different sampling ratios for comparing GI with AIME and existing encoding patterns.
Figure A11: Experimental imaging results of scene 2 under different sampling ratios for comparing GI with AIME and existing encoding patterns.
Figure A12: Experimental imaging results of scene 3 under different sampling ratios for comparing GI with AIME and existing encoding patterns.
Figure A13: Experimental imaging results of scene 4 under different sampling ratios for comparing GI with AIME and existing encoding patterns.
Appendix F Information-optimal encoding under different SNRs
Figures A14 and A15 show AIME results corresponding to the same imaging scene pepper under three different SNRs, which indicates that AIME can adaptively deal with different noise levels and constraints, resulting in different encoding patterns.
Figure A14: Information-optimal encoding from AIME-$\ell_{1}$ under three different SNRs.
Figure A15: Information-optimal encoding from AIME-$\ell_{\infty}$ under three different SNRs.